OpenStudy (anonymous):

Find all values of k so that the giver points are sqt/29 units apart. (-5,5), (k,0) Using the distance formula I get 4=(k+5)^2 What do I do next?

OpenStudy (anonymous):

(k+5)^2 +(0-5)^2 = 29 (k+5)^2=4 k+5=+-2 k=-5+-2 k=-7 or -3

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

ahh.. so yo keep the "+" and "-" in the equation and then you just simplify both. Thank you!

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

If x^2=a then